1 1
--- + --- = 1 Solve by using quadratic formula
y-5 y+5

Answer by stanbon(75887)

1 1
--- + --- = 1 Solve by using quadratic formula
y-5 y+5
-----------------------
Multiply thru by (y-5)(y+5) to get:
y+5 + y-5 = (y-5)(y+5)
2y = y^2-25
---
Rearrange:
y^2 -2y -25 = 0
---
Quadratic Formula:
y = [2 +- sqrt(2^2-4*-25)]/2
-----
y = [2 +- sqrt(104)]/2
---
y = [2 +- 2sqrt(21)]/2
---
y = 1+sqrt(21) or y = 1-sqrt(21)
====================================
Cheers,
Stan H.
==============